Isaac arrived at LML at 9:30am both doors on the fog collector had not opened and no sample had been collected. Upon closer examination the system had appeared to have been turned off prior to arrival. It was later determined that this was done during the previous day when the collector was serviced. The collector was turned back on and immediately the system began to run the program to collect the fog. the system at this point was operational however when the humidity measurement was increased by blowing on the sensor the system began to initiate the process to collect fog. Sounds for the solenoids to fire were noted but the solenoids did not fire and the system froze. The system was turned off and a test was run. The test showed that the solenoids again did not fire but the big fan did turn on for awhile until it automatically went off. A manual test was then administered to test the firing of the solenoids and the big fan. The results of this test showed that the solenoids did not fire but the big fan did fire. The system was then turned off in order to determine further causes of the crashing.
Monthly Archives: July 2014
BML 7/29
No new sample, maybe tonight
140725/26/27 LML
On the 25th Kona found that the doors were open but the system had crashed and one of the wires to the small fan was broken. She reset the system and dismantled the apparatus to wash with 409 at the sink. During the cleaning process the system crashed again. A system test ran smoothly and the system was reset.
The next morning the system had crashed again after the doors opened. Kona tested the system manually by increasing the humidity on the RH sensor. The small fan could not run due to a faulty wire, the doors opened, then there was a click sound but the big fan did not turn on and the system froze. Kona rinsed the apparatus with DI. The system was reset again in hope of something different the next day.
The morning of the 27th Kona arrived to find the same situation had occurred. The apparatus appeared clean so she closed the doors and shut down the system.
140728 UCSC
Kona arrived at 9 am, there was less that 10 mL of water collected so it was discarded. The apparatus was washed in place with 409 and a thorough DI rinse. The collecting bottle was rinsed at the sink. System test was fine, ready for fog collection.
140727 Fritzche Field
This is a report on the collection activities from the past 4 days.
Thursday 7/24: I was surprised to see a sample this morning, a little over 100 ml. When I arrived at 8:30 the doors were open and the fan was running yet the fog had ended. I rinsed the system, but closed it down from the fogPI, since I did not want it to be open all day. I knew someone would come later to take a blank and they could restart it then. Chris and Erin came at 1 pm and took a blank, cycled power, and put on a bottle.
As an aside, a spider trap was placed a day or 2 before.
Friday 7/25: When I drove by the system that morning, expecting a sample, it was closed up. Apparently, cycling the power was insufficient to restart fogPI. I came by later that afternoon and found that the bottle had also fallen off. I reinitiated the fogPI through the system and installed a new bottle.
Saturday 7/26: No sample, though we thought there had been fog. It had opened. Chris, Erin, Cristian and I acid washed the system.
Sunday 7/27: The bottle had fallen off again! This was the 3rd time! We need to tighten it fully and then just slightly loosen it to allow air to flow out, but not loosen it more than just that slight amount due to the conditions of the site. Since it was opened and bugs had flown in, I carefully cleaned the strings by pulling them out and rinsing them with DI water. With a clean hand, I reinstalled them.
No spiders yet.
140727 ISB(UCSC)
Robert arrived at ISB at 8:30. The doors were open but no sample had been collected. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.
140726 ISB(UCSC)
Robert arrived at ISB at 8:30. The doors were open but no sample had been collected. One of the strand frames came untied so only 2 frames are currently in the collector. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.
140724 UCSC Observations
Isaac arrived at the top of ISB at 9:30 am. The doors to the collector had opened but no sample had been collected. The collector was disassembled and the body was sprayed with 409 followed by a thorough rinse with DI water. The collector components were additionally sprayed with 409 followed by a thorough DI water rinse. The collector was then reassembled for the next fog event.
140725 ISB(UCSC)
Robert arrived at ISB at 7:45. The doors were open but no sample had been collected. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.
BML Update 7/25
New memory card installed to Rpi with an RH of 91. System started up and is functioning properly. Unknown if we have a sample yet (from last night), will be checking today. But system is back up and running!